# supReferences

### v4

@@ A.3.3 Proof of Theorem 1
$\mu_{t+1}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{F^{-1}_{\hat{\mu}_{t}^{N}M}(U_{t+1,1}^{n})},\quad\hat{\mu}^{N}_{t}=\Psi_{G}(\mu_{t}^{N}),\quad\forall t\geq 1.$
Then, it is direct to see that we have $(\hat{\eta}_{t}^{N})_{t\geq 1}\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{dist}}}}}{{=}}(\hat{\mu}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{dist}}}}}{{=}}(\mu_{t}^{N})_{t\geq 1}$.
Using these two results, it follows that to prove the theorem we can use the version of Lemma 1 we obtain if, in Lemma 8, we replace the right-hand size of (29) by $2D_{1}^{*}(\{u_{1}^{n}\}_{n=1}^{N})$, that is, to prove the lemma we can do as if the conclusion of Lemma 1 holds with $\delta_{N,T}=\sup_{t\in\{1,\dots,T\}}D_{1}^{*}(\{U_{t,1}^{n}\}_{n=1}^{N})^{2}$. Using this observation the result of the theorem then follows from Lemma 3. ∎
A.3.4 Proof of Proposition 3
Proof.
Let $\kappa\in(0,1)$ and let $T_{\kappa}\in\mathbb{N}$ and $a_{\kappa}\in(0,\infty)$ be such that we have both $\eta_{t}([-a_{\kappa},a_{\kappa}])\leq\kappa$ and $\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])\leq\kappa$ for all $T\geq T_{\kappa}$. Remark that such constants $T_{\kappa}$ and $a_{\kappa}$ exist by Proposition 1.
Next, let $\varrho_{\kappa}=\min\big{\{}1-\Phi(4a_{\kappa}/\sigma),1/2\big{\}}$ so that, using (5) and with $(\mathcal{F}^{N}_{t})_{t\geq 1}$ as defined in Section 3.1, we have
$\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}^{N}_{t-1}\Big{)}\geq\varrho^{N}_{\kappa},\quad\forall t\geq 1.$ (61)
Noting that for all $t\geq T_{\kappa}$ we have
$1-\mathbb{P}\Big{(}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa|\mathcal{F}_{t-1}^{N}\Big{)}\geq\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}_{t-1}^{N}\Big{)}$
it follows from (61) that
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa\Big{)}\leq(1-\varrho^{N}_{\kappa})^{T-T_{\kappa}+1}.$
To complete the proof of the proposition let $q\in(0,1)$ and note that
$(1-\varrho^{N}_{\kappa})^{T-T_{\kappa}+1}<1-q,\quad\forall N<\frac{\log\big{(}(1-(1-q)^{\frac{1}{T-T_{\kappa}+1}})^{-1}\big{)}}{\log(1/\varrho_{\kappa})}\leq\frac{\log(T-T_{\kappa}+1)-\log(q)}{\log(1/\varrho_{\kappa})}$
where the inequality uses the fact that, by Bernoulli’s inequality (see e.g. Li and Yeh,, 2013), we have $(1-q)^{1/(T-T_{\kappa}+1)}\leq 1-q/(T-T_{\kappa}+1)$. The result of the proposition follows. ∎

### v5

@@ 5.4.1 Two preliminary results
$\begin{split}\sup_{a\in\mathbb{R}}\big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi_{p}(\hat{\mu}_{p-1})\big)(f_{a})\big|&\leq\frac{\bar{C}}{\iota^{\prime\prime\prime}\iota_{T-1}}\,|\rho_{t-p}|^{\gamma}\Delta_{T}^{1-\gamma},\quad\forall\gamma\in[0,1]\end{split}$ (52)
with $\iota^{\prime\prime}=\min\{\iota^{\prime},\iota^{\prime\prime}\}>0$.
The result in (52) holds for all $t\in\{2,\dots,T\}$ and all $p\in\{1,\dots,t-1\}$ (assuming that $T\geq 3$), and we now show that it also holds when $p=t$ for some $t\in\{1,\dots,T\}$, and with $T\geq 1$ arbitrary. To do so we let $\gamma\in[0,1]$, $t\in\{1,\dots,T\}$ and $p=t$, and we note that
$\sup_{a\in\mathbb{R}}\big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\big|$ $=\sup_{a\in\mathbb{R}}\big|\hat{\mu}_{t}(f_{a})-\phi(\hat{\mu}_{t-1})(f_{a})\big|$
$=\sup_{a\in\mathbb{R}}\big|\Psi_{G}(\mu_{t})(f_{a})-\Psi_{G}(\hat{\mu}_{t-1}M)(f_{a})\big|$
$\leq\frac{5\|\mu_{t}-\hat{\mu}_{t-1}M\|}{\hat{\mu}_{t-1}M(G_{c})}$
$\leq\frac{5\Delta_{T}}{\hat{\mu}_{t-1}M(G_{c+c_{\star}})}$
Recalling that $\hat{\mu}_{k}M(G_{c+c_{\star}})\geq\iota^{\prime\prime}\iota_{T-1}\geq\iota^{\prime\prime\prime}\iota_{T-1}$ for all $k\in\{0,\dots,T-1\}$, and noting that since we are assuming that $p=t$ we have $\rho_{t-p}=\rho_{0}=1$, it follows that
$\sup_{a\in\mathbb{R}}\big|\phi^{p-t}(\hat{\mu}_{p})(f_{a})-\phi^{p-t}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\big|$ $\leq\bar{C}(\iota^{\prime\prime\prime}\iota_{T-1})^{-1}\min\big\{1,\Delta_{T}\big\}$
$\leq\bar{C}(\iota^{\prime\prime\prime}\iota_{T-1})^{-1}\min\big\{1,\Delta_{T}^{1-\gamma}\big\}$
$\leq\bar{C}(\iota^{\prime\prime\prime}\iota_{T-1})^{-1}|\rho_{t-p}|^{\gamma}\Delta_{T}^{1-\gamma}$
showing that (52) holds for any integers $1\leq p\leq t\leq T$.
We now let $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, so that $|\rho_{k}|\leq C_{\star}\epsilon_{\star}^{k}$ for all $k\geq 0$. In addition, we let $t\in\{1,\dots,T\}$ and $k\in\mathbb{N}$. Then, by using (43) and (52), and using the convention that empty sums equal zero, we have
$\begin{split}\|\hat{\mu}_{t}-\hat{\eta}_{t}\|&=\sup_{a\in\mathbb{R}}\big|\hat{\mu}_{t}(f_{a})-\hat{\eta}_{t}(f_{a})\big|\\ &\leq\sum_{p=1}^{t}\sup_{a\in\mathbb{R}}\Big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\Big|\\ &=\sum_{p=1}^{t-k}\sup_{a\in\mathbb{R}}\Big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\Big|\\ &+\sum_{p=t-k+1}^{t}\sup_{a\in\mathbb{R}}\Big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\Big|\\ &\leq\frac{\bar{C}}{\iota^{\prime\prime\prime}\iota_{T-1}}\bigg(C_{\star}\sum_{p=1}^{t-k}\epsilon_{\star}^{t-p}+k\Delta_{T}\bigg)\\ &\leq\frac{\bar{C}}{\iota^{\prime\prime\prime}\iota_{T-1}}\bigg(\epsilon_{\star}^{k}C_{\star}(1-\epsilon_{\star})^{-1}+k\Delta_{T}\bigg).\end{split}$ (53)
By applying (53) with
$k=\bigg\lceil\frac{1}{\log(1/\epsilon_{\star})}\log\bigg(\frac{C^{\prime\prime}\log(1/\epsilon_{\star})}{\Delta_{T}}\bigg)\bigg\rceil,\quad C^{\prime\prime}:=C_{\star}(1-\epsilon_{\star})^{-1},$
we obtain that
$\|\hat{\mu}_{t}-\hat{\eta}_{t}\|$ $\leq\Delta_{T}\,\bigg(\frac{\bar{C}}{\iota^{\prime\prime\prime}\iota_{T-1}\log(1/\epsilon_{\star})}\bigg)\bigg(2+\log(C^{\prime\prime})+\log\big(\log(1/\epsilon_{\star})\big)-\log(\epsilon_{\star}\Delta_{T})\bigg)$
$\leq C^{\prime\prime\prime}\Delta_{T}\log(1+\Delta_{T}^{-1})$
where the second inequality holds for some constant $C^{\prime\prime\prime}\in(0,\infty)$ depending only on $C^{\prime\prime}$ and $\epsilon_{\star}$, and uses the fact that $\Delta_{T}\in[0,1]$. The first part of the lemma follows.
To show the second part of the lemma assume first that $T\geq 2$, let $t\in\{2,\dots,T\}$, and note that
$\|\mu_{t}-\eta_{t}\|$ $\leq\|\mu_{t}-\hat{\mu}_{t-1}M\|+\|\hat{\mu}_{t-1}M-\hat{\eta}_{t-1}M\|\leq\Delta_{T}+\|\hat{\mu}_{t-1}M-\hat{\eta}_{t-1}M\|.$ (54)
Noting that for all $a\in\mathbb{R}$ we have
where, using Lemma 3, $V(\tilde{g}_{a})=V(h_{a,\rho,\sigma})\leq 1$, it follows from Lemma 4 that
$\|\hat{\mu}_{t-1}M-\hat{\eta}_{t-1}M\|=\sup_{a\in\mathbb{R}}|\hat{\mu}_{t-1}(\tilde{g}_{a})-\hat{\eta}_{t-1}(\tilde{g}_{a})|\leq\|\hat{\mu}_{t-1}-\hat{\eta}_{t-1}\|.$
By combining this latter result with (54) we obtain that
$\sup_{t\in\{2,\dots,T\}}\|\mu_{t}-\eta_{t}\|\leq\Delta_{T}+\sup_{t\in\{1,\dots,T\}}\|\hat{\mu}_{t}-\hat{\eta}_{t}\|.$
Then, the second part of the lemma follows from the first part of the lemma, upon noting that $\|\hat{\mu}_{1}-\eta_{1}\|=\|\hat{\mu}_{1}-\hat{\mu}_{0}^{N}M\|\leq\Delta_{T}$. This ends the proof of the lemma. ∎
5.4.2 Proof of the lemma
Proof of Lemma 1.
Let $(\mu_{t})_{t\geq 1}$ and $(\hat{\mu}_{t})_{t\geq 1}$ be as in the statement of the lemma, and let $I=[-\xi,\xi]$ for some $\xi\in(0,\infty)$ such that $\eta_{1}(I)>0$. Then, by using Lemma 5, we can easily verify verify that $\iota_{1}:=\inf_{t\geq 1}\hat{\eta}_{t}(I)>0$, which implies that $\iota_{2}:=\inf_{t\geq 2}\eta_{t}(I)\geq\iota_{1}\inf_{x\in I}M(x,I)>0$. Consequently, there exists a constant $\iota\in(0,1)$ such that $\inf_{t\geq 1}\eta_{t}(I)\geq\iota$.
To prove the lemma we let $\tilde{\Delta}_{T}=\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\hat{\mu}_{t-1}M\|_{\mathrm{D}}$ for all $T\geq 1$, with the convention that $\hat{\mu}_{0}M=\eta_{1}$, and we remark first that, for all $t\geq 1$,
$\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq\iota/2\implies\mu_{t}(I)=\eta_{t}(I)+\mu_{t}(I)-\eta_{t}(I)\geq\eta_{t}(I)-\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\geq\iota/2.$ (55)
Then, by using (2), (55) and Lemma 6, to prove the lemma it suffices to show that there exists an $\epsilon\in(0,1)$ such that
$\tilde{\Delta}_{T}\leq\epsilon\implies\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq\iota/2,\quad\forall T\geq 1.$ (56)
To do so we let $C_{I}\in[1,\infty)$ be as in the statement of Lemma 6 (when $I=[-\xi,\xi]$), $\delta\in(0,1)$ be arbitrary, and $C^{\prime}=\max\big\{1,\sup_{x\in(0,1]}x^{1-\delta}\log(1+1/x)\big\}<\infty$. With this notation in place, we show by induction on $T\geq 1$ that (56) holds for $\epsilon=(\iota^{2}/(8C_{I}\,C^{\prime}))^{1/\delta}$.
Since $\epsilon\leq\iota/2$ and $\tilde{\Delta}_{1}=\|\mu_{1}-\eta_{1}\|_{\mathrm{D}}$, the implication in (56) holds when $T=1$, and we now assume that it holds when $T=s$ for some $s\geq 1$. Then, by using (55), it follows that $\mu_{t}(I)\geq\iota/2$ for all $t\in\{1,\dots,s\}$ and thus, by Lemma 6 and using (2), we have
$\sup_{t\in\{1,\dots,s+1\}}\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq 2\sup_{t\in\{1,\dots,s+1\}}\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq\frac{4C_{I}}{\iota}\Delta_{s+1}\log(1+\Delta_{s+1}^{-1})$ (57)
where
$\frac{4C_{I}}{\iota}\Delta_{s+1}\log(1+\Delta_{s+1}^{-1})\leq\frac{\iota}{2}\Leftrightarrow\Delta_{s+1}\log(1+\Delta_{s+1}^{-1})\leq\frac{\iota^{2}}{8C_{I}}.$ (58)
By the definition of the constant $C^{\prime}$, we have $\Delta_{s+1}\log(1+\Delta_{s+1}^{-1})\leq C^{\prime}\,\Delta_{s+1}^{\delta}\leq C^{\prime}\,\tilde{\Delta}_{s+1}^{\delta}$ which, together with (57) and (58), shows that $\tilde{\Delta}_{s+1}\leq\epsilon\implies\sup_{t\in\{1,\dots,s+1\}}\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq\iota/2$, showing that (56) holds for $T=s+1$. This concludes the proof of (56) and thus of the lemma.